OpenStudy (anonymous):

12/17 divided by 4/17 = ? 4 48 3 17

OpenStudy (saifoo.khan):

\[\frac{12}{17} \div \frac{4}{17}\] \[\frac{12}{17} \times \frac{17}{4}\] \[\frac{12}{\cancel{17}} \times \frac{\cancel{17}}{4}\]

mathslover (mathslover):